Florida Hospital Apopka, a not-for profit organization, began 2018 with the following account balances on January 1:

Cash

$70,000

Accounts receivable

245,000

Allowance for doubtful accounts

18,000

Supplies inventory

24,000

Property and Equipment

1,500,000

Accumulated depreciation

300,000

Accounts payable

21,000

Notes payable (short-term bank loans)

500,000

Net assets

1,000,000

During 2018, the accounting clerk recorded the following transactions (Florida Hospital Apopka’s year end is December 31):

Transaction Number

Event

Amount

1.

Billed patients for services rendered

$1,700,000

2.

Purchased medical supplies on credit

12,000

3.

Employee salaries earned during 2018

712,000

4.

Employee salaries paid during 2018

683,000

5.

Annual depreciation on Property and Equipment

150,000

6.

Received a bank loan (short-term)

750,000

7.

Cash collections on patient billings

1,124,000

8.

Estimated bad debts for 2018 billings

44,000

9.

Made payment on bank loan

575,000

10.

Used medical supplies inpatient care

10,000

a. Prepare a properly formatted 2018 balance sheet and income statement for the hospital using the beginning account balances and incorporating the effects of each transaction.

b. What is Florida Hospital Apopka’s net working capital for 2018?

c. What is Florida Hospital Apopka’s debt ratio?

d. How does Florida Hospital Apopka’s debt ratio compare with the debt ratio for Sunnyvale (textbook, Exhibit 4.1)?
Answer #1
a) INCOME STATEMENT
Service revenue 1700000
Operating expenses:
Salaries expense 712000
Supplies expense 10000
Depreciation expense-PPE 150000
Bad debts 44000 916000
Net income 784000
BALANCE SHEET
Current Assets:
Cash (70000-683000+750000+1124000-575000) 686000
Accounts receivables (245000+1700000-1124000) 821000
Allowance for doubtful accounts (18000+44000) -62000
Medical supplies (24000+12000-10000) 26000 1471000
Non-Current Assets:
Property and equipment 1500000
Accumulated depreciation (300000+150000) 450000 1050000
Total assets 2521000
Current liabilities:
Accounts payable (21000+12000) 33000
Salaries payable (712000-683000) 29000
Notes payable (500000+750000-575000) 675000 737000
Equity (1000000+784000) 1784000
Total liabilities & Equity 2521000
b) Net working capital = 1471000-737000 = 734000
c) Debt ratio = 737000/2521000 = 29.23%
d) Debt ratio of Sunnyvale is required to answer this
part.